A coin is tossed, then a letter from the word INDIANAPOLIS is selected at random. Find each probability. P( heads, then P) P( tails, then a vowel) P( tails, then N) P( heads, then D or I)


1
Expert's answer
2020-11-12T18:40:59-0500

"Pr(Head)=\dfrac{1}{2}=P(Tail)"

There are 12 letters in the word INDIANAPOLIS: 6 vowels and 6 consonants.

"Pr(P)=\dfrac{1}{12}"

"Pr(Vowel)=\dfrac{6}{12}=\dfrac{1}{2}"

"Pr(N)=\dfrac{2}{12}=\dfrac{1}{6}"

"Pr(D\cup I)=\dfrac{4}{12}=\dfrac{1}{3}"


Then


"Pr(Head, P)=\dfrac{1}{2}\cdot \dfrac{1}{12}=\dfrac{1}{24}"

"Pr(Tail, Vowel)=\dfrac{1}{2}\cdot \dfrac{1}{2}=\dfrac{1}{4}"

"Pr(Tail, N)=\dfrac{1}{2}\cdot \dfrac{1}{6}=\dfrac{1}{12}"

"Pr(Head, D\cup I)=\dfrac{1}{2}\cdot \dfrac{1}{3}=\dfrac{1}{6}"